OpenStudy (anonymous):

x + y = k x - y = k The solution to the system shown is _____. (0, k) (2, 2) (k, 0)

OpenStudy (anonymous):

Plug in the given values and see if it works for both equations. Can you give it a try?

OpenStudy (anonymous):

k+0=k k-0=k (k,0)